OpenStudy (anonymous):

solve 63(4^[4-3x])=7

OpenStudy (anonymous):

would it be x=2/3 ?

OpenStudy (anonymous):

no it x=2.41\3

OpenStudy (anonymous):

how so

OpenStudy (anonymous):

first div 63 on both sides 4^[4-3x] = 1/9 take log on both sides

OpenStudy (anonymous):

oh okay. so how would I solve 5^[x+2] + 5^x = 24

OpenStudy (anonymous):

when the value of x is in power of any coeff then to try to use log method

OpenStudy (anonymous):

um okay, but can't I take the common value? which is 5^x?

OpenStudy (anonymous):

take the common factor 5^x(5^2+1)=24 5^x(25+1)=24 5^x(26)=24